# Measuring Diversity in Synthetic Datasets

### v1

@@ Appendix A Additional Related Work
Limited by the space, we provide a literature review of the LLM dataset generator and application of diversity evaluation methods as follows.
@@ A.1 LLM Dataset Generator
Prompt-guided and Dataset-guided Strategies. Recent studies (Ding et al., 2022; Chung et al., 2023) leverage LLMs to augment existing datasets or generate a dataset from scratch, demonstrating the effectiveness in improving dataset quality and reducing data collection costs. Generally, efforts to employ LLMs as dataset generators can be categorized into three strategies: Prompt-guided (Li et al., 2023), Dataset-guided (Ye et al., 2022), and Instruct-guided (Samuel et al., 2023). The prompt-guided strategy, a prominent data augmentation approach using LLMs, involves designing task-specific prompts to guide LLMs to augment data in a few-shot (Yoo et al., 2021) or zero-shot (Mahmoudi et al., 2024) manner. Due to its simplicity and effectiveness, subsequent works extend this strategy to various scenarios, such as medical (Yuan et al., 2023), person retrieval (Li et al., 2024b), and social media scenario (dos Santos et al., 2024). However, simple prompt engineering has limitations in fully exploiting the capabilities of LLMs, leading to the development of multi-level prompt designs (Ye et al., 2024) and targeted sample augmentation (Yang et al., 2024b). To further harness the potential of LLMs, the dataset-guided strategy employs LLMs to generate a training set and then trains a task-specific model to annotate unlabeled data (Sahu & Laradji, 2024). The dataset-guided strategy aims to approximate the distribution of targeted scenarios, but it is currently only applicable to text classification tasks.
Instruct-guided Strategy. Previous studies (White et al., 2023) indicate that the design of prompts significantly impacts the performance of LLMs, spurring research into the instruct-guided strategy. Generally speaking, the instruct-guided strategy leverages LLMs to generate instructions that guide another LLM in dataset generation (Evuru et al., 2024). These instructions typically relate to context (Samuel et al., 2023), criteria (Huang et al., 2023), and tasks (Wang et al., 2022). To further improve the quality of instructions, efforts have been concentrated on selecting optimal instructions (Li et al., 2024a), integrating soft instructions (Chen et al., 2023), and implementing self-correction mechanisms (Gupta et al., 2023). In a nutshell, LLMs are employed to generate or augment datasets through prompt engineering and multi-step strategies, which encompass various application scenarios and downstream tasks. Meanwhile, the diversity of synthetic datasets emerges as a critical factor in measuring data quality. In our work, we focus on the diversity evaluation of synthetic datasets derived from any dataset generation strategies.
@@ A.2 Application of Diversity Evaluation Methods
Quantifying Augmentation Performance. As data augmentation becomes an essential component in the training of deep neural networks (Zhang, 2017; Park et al., 2019), researchers gradually explore a better quantification of the quality of data augmentation. Some studies (Cubuk et al., 2019) suggest that the effectiveness of data augmentation arises from the increased diversity of the data. Inspired by this observation, a series of studies have introduced diversity evaluation metrics into the performance assessment of data augmentation strategies. Specifically, they consider diversity as one aspect of evaluating the quality of augmented data, thereby determining the effectiveness of data augmentation. For instance, (Gontijo-Lopes et al., 2020) utilize the fundamental idea that models find it more challenging to fit more diverse data, comparing metrics such as training loss and training time before and after augmentation to assess diversity. Similarly, (Yang et al., 2024a) evaluate diversity by examining the eigenvalues and eigenvectors of the similarity matrix of samples before and after augmentation.
@@
Other Applications. In addition to the aforementioned applications, diversity evaluation metrics have valuable applications in various areas, including sample selection for datasets (Cao et al., 2023), enhancing robustness (Lee et al., 2022), and eliminating biases within datasets (Huber et al., 2024).
@@ Appendix B Proof of Properties of DCScore
We theoretically confirm that DCScore satisfies several intuitive axioms pointed out by previous studies (Leinster & Cobbold, 2012), thereby demonstrating its superiority.
@@
    Assuming that $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are completely identical, and the samples within each dataset are entirely different, i.e., $\operatorname{DCScore}(\mathcal{D}_{1})=\operatorname{DCScore}(\mathcal{D}_{2})=n$. Let $\textbf{P}=[P_{1},...,P_{n},...,P_{2n}]$ denote the probability matrix of the merged dataset $\mathcal{D}^{{}^{\prime}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}=\{\mathcal{T}_{i}\}_{i=1}^{2n}$. For $1\leq i\leq n$, $\mathcal{T}_{i}=\mathcal{T}_{2i}$, where $\mathcal{T}_{i}\in\mathcal{D}_{1}$, $\mathcal{T}_{2i}\in\mathcal{D}_{2}$. Consequently, for each diversity-sensitive component $\mathcal{\tilde{T}}_{i}$ in $\mathcal{D}^{{}^{\prime}}$, $\textbf{P}[i,i]=\textbf{P}[i,2i]=\frac{1}{2}$. Finally, $\operatorname{DCScore}(\mathcal{D}^{{}^{\prime}})=\sum_{i=1}^{2n}{\frac{1}{2}}=n$.
@@
    According to (4), for the $i$-th diversity-sensitive component in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, the probability of being classified as category $c_{i}$ can be computed as follows:
@@
    For a merged dataset $\mathcal{D}^{{}^{\prime}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}=\{\mathcal{T}_{i}\}_{i=1}^{2n}$, when $1\leq i\leq n$, we have $\mathcal{T}_{i}=\mathcal{T}_{2i}$, where $\mathcal{T}_{i}\in\mathcal{D}_{1}$, $\mathcal{T}_{2i}\in\mathcal{D}_{2}$. Since the newly added data samples do not affect the pairwise similarity, the kernel matrix $\textbf{K}^{{}^{\prime}}$ for $\mathcal{D}^{{}^{\prime}}$ can be formulated as follows:
@@
    Analogous to (11), for $1\leq i\leq n$, the probability of the $i$-th diversity-sensitive component in $\mathcal{D}^{{}^{\prime}}$ being classified as category $c_{i}$ can be computed as follows:
@@
    For $n+1\leq i\leq 2n$, we obtain the same result as depicted in (13). Consequently, the diversity of $\mathcal{D}^{{}^{\prime}}$ can be computed as follows:
@@
  • Symmetry (Restated): Diversity remains constant regardless of the order of the samples, exhibiting permutation invariance. Let $\pi$ denote the permutation function for the sample order, DCScore remains unchanged for any sample permutation of $\mathcal{D}$, i.e.,
@@
    According to (4), the order of samples does not affect the classification task. Thus, the diagonal elements of P remain unchanged, indicating the symmetry property of DCScore. ∎
  • Monotonicity (Restated): The diversity of a dataset decreases as the similarity between its samples increases. Given two datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and a new sample $\mathcal{T}_{n+1}$, where the samples in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are entirely different, and $\operatorname{DCScore}(\mathcal{D}_{1})=\operatorname{DCScore}(\mathcal{D}_{2})=n$. If $\mathcal{T}_{n+1}$ is more similar to the samples in $\mathcal{D}_{2}$ than to those in $\mathcal{D}_{1}$ and is added to both datasets, then for the merged datasets $\mathcal{D}_{1}^{{}^{\prime}}$ and $\mathcal{D}_{2}^{{}^{\prime}}$, DCScore satisfies the following equation.
@@
    For $\mathcal{D}_{1}^{{}^{\prime}}=\{\mathcal{T}_{1}^{1},\mathcal{T}_{2}^{1},...,\mathcal{T}_{n}^{1},\mathcal{T}_{n+1}\}$ and $\mathcal{D}_{2}^{{}^{\prime}}=\{\mathcal{T}_{1}^{2},\mathcal{T}_{2}^{2},...,\mathcal{T}_{n}^{2},\mathcal{T}_{n+1}\}$, we have $\operatorname{S}(\mathcal{T}_{i}^{1},\mathcal{T}_{n+1})<\operatorname{S}(\mathcal{T}_{j}^{2},\mathcal{T}_{n+1})$ for any $i,j=\{1,2,...,n\}$. Here, $\operatorname{S}$ is the similarity function. In this regard, the classification function $f$ exhibits lower confidence when classifying dataset $\mathcal{D}_{2}^{{}^{\prime}}$, resulting in a lower probability that the $i$-th sample is classified into the $i$-th class, thereby leading to $\textbf{P}_{\mathcal{D}_{1}^{{}^{\prime}}}[i,i]>\textbf{P}_{\mathcal{D}_{2}^{{}^{\prime}}}[i,i]$. Then, the following formula is satisfied:
@@
Appendix C Algorithm and Other Implementations for a Classification Perspective
C.1 Other Implementations
The process in Section 4.1 represents one way to implement the classification perspective of DCScore. There are other potential implementations of DCScore with lower complexity. When $n\gg d$, the classification probability modeling for any sample within the evaluated dataset is primarily determined by samples that are relatively similar to it, while the majority of other samples fall into the tail probability category. Based on this observation, we can reduce computational cost by calculating pairwise similarity only among similar samples. Therefore, we propose a feasible approach named $\texttt{DCScore}_{cluster}$ below.
$\texttt{DCScore}_{cluster}$ defines similar samples as those within the same cluster. Specifically, $\texttt{DCScore}_{cluster}$ involves obtaining cluster centers and the membership of each sample point through clustering, followed by calculating pairwise similarities among samples within the same cluster. By performing clustering, we identify similar samples for each sample point to facilitate pairwise similarity calculation, thereby avoiding the computation of global pairwise similarities and further reducing the computational complexity of the method. Based on this process, the method can be formulated as follows:
Text Representation: H=Φ($\begin{split}&\textnormal{{Text Representation: }}\textbf{H}=\Phi(\{\mathcal{\tilde{T}}_{i}\}_{i=1}^{n}),\\ &\begin{aligned} \textnormal{{Pairwise Similarity: }}&\mathcal{C}=\operatorname{K-means}(\textbf{H}),\\ &\textbf{K}[i,j]=\begin{cases}\operatorname{Kernel}(\textbf{h}_{i},\textbf{h}_{j}),&\text{if }\textbf{h}_{i},\textbf{h}_{j}\in\textbf{C}_{k}\\ 0,&\text{if }\textbf{h}_{i}\in\textbf{C}_{k}\text{ and }\textbf{h}_{j}\in\textbf{C}_{l}\text{ with }k\neq l\end{cases}\end{aligned}\\ &\begin{aligned} \textnormal{{Diversity Summarization: }}&\textbf{P}[i,j]=\begin{cases}f_{\textbf{K}}(\textbf{K}[i,j]),&\text{if }\textbf{h}_{i},\textbf{h}_{j}\in\textbf{C}_{k}\\ 0,&\text{if }\textbf{h}_{i}\in\textbf{C}_{k}\text{ and }\textbf{h}_{j}\in\textbf{C}_{l}\text{ with }k\neq l\end{cases}\\ &\operatorname{DCScore}(\mathcal{D})=\operatorname{tr}(\textbf{P})=\sum_{i=1}^{n}{\textbf{P}[i,i]}.\end{aligned}\end{split}$ (18)
where $\textbf{C}_{k},\textbf{C}_{l}\in\mathcal{C}$ are clusters.
Appendix D Experimental Settings
D.1 Datasets
@@
D.1.1 Self-generated Datasets
@@
We utilize two different self-generated datasets in three subsections: Section 5.2.1, Section 5.2.2, and Section 5.4. It is worth noting that Section 5.2.1 and Section 5.4 share the same experimental datasets. We employ two commonly used LLMs as our dataset generator, including Llama2-13B (13B) and Llama2-70B (70B) (Touvron et al., 2023). To prompt LLMs to generate datasets, we design two prompts corresponding to Zero-shot and Few-shot generation settings, respectively. Additionally, self-generated datasets involve two natural language generation (NLG) tasks: text classification and story completion. We set the maximum number of generated tokens to 100 and 30 for text classification and story completion tasks, respectively. The detailed generation information is offered as follows.
Generation Settings. We use different generation settings for generated datasets used in Section 5.2.1, Section 5.2.2, and Section 5.4.
@@
  • Datasets on Section 5.2.1, Section 5.4, Appendix E.3, and Appendix E.4. We generate 21 sub-datasets corresponding to different $\tau_{g}$ by varying $\tau_{g}$ from 0.2 to 1.2 with 0.05 intervals. For each sub-dataset, we employ LLMs (13B or 70B) to generate sets of 10 responses per context. Specifically, each sub-dataset consists of 100 samples.
@@
  • Datasets on Section 5.2.2. We employ the 70B model to generate 6 sub-datasets corresponding to different $\tau_{g}$ by varying $\tau_{g}$ from 0.2 to 1.2 with 0.2 intervals. Each sub-dataset includes 5 samples corresponding to a context. To repeat experiments five times, we use five different contexts to prompt the 70B model to generate 5 sub-datasets for each $\tau_{g}$.
@@
  • Datasets on Appendix 5.5. In zero-shot or few-shot settings, we utilize the 70B model to generate three sub-datasets for the text classification task, corresponding to $\tau_{g}=\{0.2,0.7,1.2\}$, respectively. Unlike other settings that provide only one example, in this experiment, we adopt a few-shot setting where four examples and their corresponding labels are given, including two positive examples and two negative examples. Each sub-dataset contains 3,000 samples, and a context is employed to prompt the 70B model to generate five samples. To train text classification models on each sub-dataset, we randomly split 2,100 samples to the training set for each sub-dataset and gather the remaining 900 samples into the testing set across all three sub-datasets. Consequently, we construct a test set comprising 1,800 samples.
@@
Prompt Settings. Following the setting of (Li et al., 2023), we design different prompts for zero-shot and few-shot settings, respectively. For the text classification task under the zero-shot setting, we require LLMs to generate movie reviews with Sci-fi/Action/Drama/Comedy/Romance topics. Each movie review contains a single sentiment of either positive or negative, which is regarded as the text label. For the story completion task, we require LLMs to complete the story according to the given context. The detailed prompt setting is provided in Table 8.
@@
D.1.2 Publicly Available Generated Datasets
We use SST2 (Socher et al., 2013), Yelp (Zhang et al., 2015), and AG News (Zhang et al., 2015), and their augmented version based on AttrPrompt (Yu et al., 2024). For three original datasets, we randomly sample data from training sets and apply this to the computational cost analysis in Section 5.3 and Appendix E.1. For three augmented datasets, each dataset has 6000 samples. We sample different sub-datasets based on these three datasets, applied to Section 5.2.1 and Section 5.3, respectively. The details are as follows.
  • Datasets on Section 5.3. We remove samples with text token lengths less than 50 in the three datasets and then truncate each sample to a length of 50 tokens. Based on the above, we set up sub-datasets with randomly sampled samples of 100, 500, 1000, 2000, and 4000 according to experimental settings.
D.2 Implementation details
For three transformation-based methods, including DCScore, VendiScore, and K-means Inertia, we employ unsup-simcse-bert-base-uncased (princeton nlp, 2021) as the embedding function. For all dataset generation language models, we set the top-p and top-k parameters to 1 and -1, respectively. Additionally, we limit the maximum number of newly generated tokens to 100 for the text classification task and 30 for the story completion task. All experiments are conducted on 8$\times$ NVIDIA Tesla V100 GPU with 32GB memory.
D.2.1 Hyperparameter Settings of Diversity Evaluation

### v2

@@ Appendix A Additional Related Work
Limited by space, we provide a literature review of the LLM dataset generator and application of diversity evaluation methods as follows.
@@ A.1 LLM Dataset Generator
Recent studies (Ding et al., 2022; Chung et al., 2023) leverage LLMs to augment existing datasets or generate a dataset from scratch, demonstrating the effectiveness in improving dataset quality and reducing data collection costs. Generally, efforts to employ LLMs as dataset generators can be categorized into three strategies: Prompt-guided (Li et al., 2023), Dataset-guided (Ye et al., 2022), and Instruct-guided (Samuel et al., 2023).
Prompt-guided and Dataset-guided Strategies. The prompt-guided strategy, a prominent data augmentation approach using LLMs, involves designing task-specific prompts to guide LLMs to augment data in a few-shot (Yoo et al., 2021) or zero-shot (Mahmoudi et al., 2024) manner. Due to its simplicity and effectiveness, subsequent works extend this strategy to various scenarios, such as medical (Yuan et al., 2023), person retrieval (Li et al., 2024b), and social media scenario (dos Santos et al., 2024). However, simple prompt engineering has limitations in fully exploiting the capabilities of LLMs, leading to the development of multi-level prompt designs (Ye et al., 2024) and targeted sample augmentation (Yang et al., 2024b). To further harness the potential of LLMs, the dataset-guided strategy employs LLMs to generate a training set and then trains a task-specific model to annotate unlabeled data (Sahu & Laradji, 2024). The dataset-guided strategy aims to approximate the distribution of targeted scenarios, but it is currently only applicable to text classification tasks.
Instruct-guided Strategy. Previous studies (White et al., 2023) indicate that the design of prompts significantly impacts the performance of LLMs, spurring research into the instruct-guided strategy. Generally speaking, the instruct-guided strategy leverages LLMs to generate instructions that guide another LLM in dataset generation (Evuru et al., 2024). These instructions typically relate to context (Samuel et al., 2023), criteria (Huang et al., 2023), and tasks (Wang et al., 2022). To further improve the quality of instructions, efforts have been concentrated on selecting optimal instructions (Li et al., 2024a), integrating soft instructions (Chen et al., 2023), and implementing self-correction mechanisms (Gupta et al., 2023).
In a nutshell, LLMs are employed to generate or augment datasets through prompt engineering and multi-step strategies, which encompass various application scenarios and downstream tasks. Meanwhile, the diversity of synthetic datasets emerges as a critical factor in measuring data quality. In our work, we focus on the diversity evaluation of synthetic datasets derived from any dataset generation strategies.
@@ A.2 Application of Diversity Evaluation Methods
Extending beyond diversity quantification, these methods demonstrate wider utility, such as quantifying augmentation performance and evaluating mode collapse. Thus, we present a literature review of other applications of the diversity evaluation.
Quantifying Augmentation Performance. As data augmentation becomes an essential component in the training of deep neural networks (Zhang, 2017; Park et al., 2019), researchers gradually explore a better quantification of the quality of data augmentation. Some studies (Cubuk et al., 2019) suggest that the effectiveness of data augmentation arises from the increased diversity of the data. Inspired by this observation, a series of studies have introduced diversity evaluation metrics into the performance assessment of data augmentation strategies. Specifically, they consider diversity as one aspect of evaluating the quality of augmented data, thereby determining the effectiveness of data augmentation. For instance, (Gontijo-Lopes et al., 2020) utilizes the fundamental idea that models find it more challenging to fit more diverse data, comparing metrics such as training loss and training time before and after augmentation to assess diversity. Similarly, (Yang et al., 2024a) evaluates diversity by examining the eigenvalues and eigenvectors of the similarity matrix of samples before and after augmentation.
@@
Other Applications. In addition to the aforementioned applications, diversity evaluation metrics have valuable applications in various areas, including sample selection for datasets (Cao et al., 2023), enhancing adversarial robustness (Lee et al., 2022) and out of distribution robustness (Leng et al., 2025), and eliminating biases within datasets (Huber et al., 2024).
@@ Appendix B Proof of Properties of DCScore
We theoretically confirm that DCScore satisfies several intuitive axioms pointed out by previous studies (Leinster & Cobbold, 2012), thereby demonstrating its role as a principled diversity evaluation method.
@@
    Assuming that $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are completely identical, and the samples within each dataset are entirely different, i.e., $\operatorname{DCScore}(\mathcal{D}_{1})=\operatorname{DCScore}(\mathcal{D}_{2})=n$. Let $\textbf{P}=[P_{1},...,P_{n},...,P_{2n}]$ denote the probability matrix of the merged dataset $\mathcal{D}^{{}^{\prime}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}=\{\mathcal{T}_{i}\}_{i=1}^{2n}$. For $1\leq i\leq n$, $\mathcal{T}_{i}=\mathcal{T}_{n+i}$, where $\mathcal{T}_{i}\in\mathcal{D}_{1}$, $\mathcal{T}_{n+i}\in\mathcal{D}_{2}$. Consequently, for each diversity-sensitive component $\mathcal{\tilde{T}}_{i}$ in $\mathcal{D}^{{}^{\prime}}$, $\textbf{P}[i,i]=\textbf{P}[i,n+i]=\frac{1}{2}$. Finally, $\operatorname{DCScore}(\mathcal{D}^{{}^{\prime}})=\sum_{i=1}^{2n}{\frac{1}{2}}=n$.
@@
    According to Eq. (4), for the $i$-th diversity-sensitive component in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, the probability of being classified as category $c_{i}$ can be computed as follows:
@@
    For a merged dataset $\mathcal{D}^{{}^{\prime}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}=\{\mathcal{T}_{i}\}_{i=1}^{2n}$, when $1\leq i\leq n$, we have $\mathcal{T}_{i}=\mathcal{T}_{n+i}$, where $\mathcal{T}_{i}\in\mathcal{D}_{1}$, $\mathcal{T}_{n+i}\in\mathcal{D}_{2}$. Since the newly added data samples do not affect the pairwise similarity, the kernel matrix $\textbf{K}^{{}^{\prime}}$ for $\mathcal{D}^{{}^{\prime}}$ can be formulated as follows:
@@
    Analogous to Eq. (11), for $1\leq i\leq n$, the probability of the $i$-th diversity-sensitive component in $\mathcal{D}^{{}^{\prime}}$ being classified as category $c_{i}$ can be computed as follows:
@@
    For $n+1\leq i\leq 2n$, we obtain the same result as depicted in Eq. (13). Consequently, the diversity of $\mathcal{D}^{{}^{\prime}}$ can be computed as follows:
@@
  • Symmetry (Restated): Diversity remains constant regardless of the order of the samples, exhibiting permutation invariance. Let $\pi(\cdot)$ denote the permutation function for the sample order, DCScore remains unchanged for any sample permutation of $\mathcal{D}$, i.e.,
@@
    According to Eq. (4), the order of samples does not affect the classification task. Thus, the diagonal elements of P remain unchanged, indicating the symmetry property of DCScore. ∎
  • Monotonicity (Restated): The diversity of a dataset decreases as the sample similarity increases. Given two datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and a new sample $\mathcal{T}_{n+1}$, where the samples in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are entirely different, and $\operatorname{DCScore}(\mathcal{D}_{1})=\operatorname{DCScore}(\mathcal{D}_{2})=n$. If $\mathcal{T}_{n+1}$ is more similar to the samples in $\mathcal{D}_{2}$ than to those in $\mathcal{D}_{1}$ and is added to both datasets, then for the merged datasets $\mathcal{D}_{1}^{{}^{\prime}}$ and $\mathcal{D}_{2}^{{}^{\prime}}$, DCScore satisfies the following equation.
@@
    For $\mathcal{D}_{1}^{{}^{\prime}}=\{\mathcal{T}_{1}^{1},\mathcal{T}_{2}^{1},...,\mathcal{T}_{n}^{1},\mathcal{T}_{n+1}\}$ and $\mathcal{D}_{2}^{{}^{\prime}}=\{\mathcal{T}_{1}^{2},\mathcal{T}_{2}^{2},...,\mathcal{T}_{n}^{2},\mathcal{T}_{n+1}\}$, we have $\operatorname{S}(\mathcal{T}_{i}^{1},\mathcal{T}_{n+1})<\operatorname{S}(\mathcal{T}_{j}^{2},\mathcal{T}_{n+1})$ for any $i,j=\{1,2,...,n\}$. Here, $\operatorname{S}(\cdot,\cdot)$ is the similarity function. In this regard, the classification function $f(\cdot)$ exhibits lower confidence when classifying dataset $\mathcal{D}_{2}^{{}^{\prime}}$, resulting in a lower probability that the $i$-th sample is classified into the $i$-th class, thereby leading to $\textbf{P}_{\mathcal{D}_{1}^{{}^{\prime}}}[i,i]>\textbf{P}_{\mathcal{D}_{2}^{{}^{\prime}}}[i,i]$. Then, the following formula is satisfied:
@@
Appendix C Experimental Settings
C.1 Datasets
@@
C.1.1 Self-generated Datasets
@@
In our experiments, we utilize three different self-generated datasets. We employ two commonly used LLMs as our dataset generator, including Llama2-13B (13B) and Llama2-70B (70B) (Touvron et al., 2023). To prompt LLMs to generate datasets, we design two prompts corresponding to Zero-shot and Few-shot generation settings, respectively. Additionally, self-generated datasets involve two natural language processing tasks: text classification and story completion. We set the maximum number of generated tokens to 100 and 30 for text classification and story completion tasks, respectively. The detailed generation information is offered as follows.
Generation Settings. We use three different generation settings across different experiments. The detailed information is shown as follows.
@@
  • Datasets on Section 5.2.1, Section 5.4, Appendix D.4, and Appendix D.5. We generate 21 sub-datasets corresponding to different $\tau_{g}$ by varying $\tau_{g}$ from 0.2 to 1.2 with 0.05 intervals. For each sub-dataset, we employ LLMs (13B or 70B) to generate sets of 10 responses per context. Specifically, each sub-dataset consists of 100 samples.
@@
  • Datasets on Section 5.2.2 and Section 5.2.3. We employ the 70B model to generate 6 sub-datasets corresponding to different $\tau_{g}$ by varying $\tau_{g}$ from 0.2 to 1.2 with 0.2 intervals. Each sub-dataset includes 5 samples corresponding to a context. To repeat experiments five times, we use five different contexts to prompt the 70B model to generate 5 sub-datasets for each $\tau_{g}$.
@@
  • Datasets on Appendix 5.5.1 and Appendix D.3. In zero-shot or few-shot settings, we utilize the 70B model to generate three sub-datasets for the text classification task, corresponding to $\tau_{g}=\{0.2,0.7,1.2\}$, respectively. Unlike other settings that provide only one example, in this experiment, we adopt a few-shot setting where four examples and their corresponding labels are given, including two positive examples and two negative examples. Each sub-dataset contains 3,000 samples, and a context is employed to prompt the 70B model to generate five samples. To train text classification models on each sub-dataset, we randomly split 2,100 samples to the training set for each sub-dataset and gather the remaining 900 samples into the testing set across all three sub-datasets. Consequently, we construct a test set comprising 1,800 samples.
@@
Prompt Settings. Following the setting of (Li et al., 2023), we design different prompts for zero-shot and few-shot settings, respectively. Here, we provide a seed example for the few-shot setting, whereas the zero-shot setting does not receive any. For the text classification task under the zero-shot setting, we require LLMs to generate movie reviews with Sci-fi/Action/Drama/Comedy/Romance topics. Each movie review contains a single sentiment of either positive or negative, which is regarded as the text label. For the story completion task, we require LLMs to complete the story according to the given context. The detailed prompt setting is provided in Table 9.
@@
C.1.2 Publicly Available Generated Datasets
We use SST2 (Socher et al., 2013), Yelp (Zhang et al., 2015), and AG News (Zhang et al., 2015), and their augmented version based on AttrPrompt (Yu et al., 2024). For three original datasets, we randomly sample data from training sets and apply this to the computational cost analysis in Section 5.3, Appendix D.1, and Appendix D.2. For three augmented datasets, each dataset has 6000 samples. We sample different sub-datasets based on these three datasets, applied to Section 5.3 and Section 5.5.3, respectively. The details are as follows.
  • Datasets on Section 5.3. We remove samples with text token lengths less than 50 in the three datasets and then truncate each sample to a length of 50 tokens. Based on the above, we set up sub-datasets with randomly selected samples of 100, 500, 1000, 2000, and 4000.
C.2 Implementation Details
For three transformation-based methods, including DCScore, VendiScore, and K-means Inertia, we employ unsup-simcse-bert-base-uncased (princeton nlp, 2021) as the weight of the embedding function. For all language models used to generate the dataset, we set the top-p and top-k parameters to 1 and -1, respectively. Additionally, we limit the maximum number of newly generated tokens to 100 for the text classification task and 30 for the story completion task. All experiments are conducted on 8$\times$ NVIDIA Tesla V100 GPU with 32GB of memory. For self-generated datasets, we only evaluate the diversity of generated components.
C.2.1 Hyperparameter Settings of Diversity Evaluation